# Optimal Convergence Rate for Mirror Descent Methods with Time-Varying Step Sizes Rules

### 09 Jan 2024 v1

@@ Abstract.
In this paper, the optimal convergence rate $O\left(N^{-1/2}\right)$ (where $N$ is the total number of iterations performed by the algorithm), without the presence of a logarithmic factor, is proved for mirror descent algorithms with time-varying step sizes, for solving both classical constrained non-smooth problems and problems with the composite model. The proven result is an improvement on the well-known rate $O\left(\log(N)N^{-1/2}\right)$ for the mirror descent algorithms with time-varying step sizes.
@@
Convex Optimization, non-smooth problem, composite problem, mirror descent, mirror-c descent, optimal convergence rate, sub-optimal convergence rate, time-varying step size.
@@ 1. Introduction
The mirror descent method which originated in [3, 4] and was later analyzed in [5], is considered as the non-Euclidean extension of standard subgradient methods. This method is used in many applications, see [10, 11, 12, 13, 14] and references therein. The standard subgradient methods employ the Euclidean distance function with a suitable step-size in the projection step. Mirror descent extends the standard projected subgradient methods by employing a nonlinear distance function with an optimal step-size in the nonlinear projection step [15]. The mirror descent method is also applicable to optimization problems in Banach spaces where gradient descent is not [16]. An extension of the mirror descent method for constrained problems was proposed in [17, 4]. Usually, the step size and stopping criterion for the mirror descent method require knowing the Lipschitz constant of the objective function and constraints, if any. Adaptive step sizes, which do not require this information, are considered for unconstrained problems in [18], and for constrained problems in [17].
As we pointed out the mirror descent algorithm was proposed and analyzed a long time ago for different step size rules. However, the optimal convergence rate (namely $O(N^{-1/2})$, where $N$ is the total number of iterations performed by the algorithm) of the algorithm was proved only with the fixed (constant) step size. When the step sizes are time-varying (non-constant in all iterations) it proved the sub-optimal convergence rate, namely $O(\log(N)N^{-1/2})$, see [1, 2, 8, 7] and references therein.
@@
The constant step size rule is relatively easy to analyze but has the disadvantage of requiring the a priori knowledge of the total number of iterations employed by the method. In practical situations, the number of iterations is not fixed a priori, and stopping criteria different than merely fixing the total number of iterations is usually imposed. This is why dynamic, or time-varying (i.e., non-constant) step size rules are important [7].
Recently in [9], it was proved the optimal convergence rate for the classical projected subgradient method for time-varying step size. This convergence rate remains the same (optimal) even if we slightly increase the weight of the most recent points, thereby relaxing the ergodic sense.
@@
In this paper, we extend the results of [9] to the mirror descent methods. We proved the optimal convergence rate, without the presence of a logarithmic factor, for mirror descent methods with time-varying step sizes, for solving both classical constrained non-smooth problems (see Sect. 3) and composite problems (see subsection 4).
@@
The paper consists of an introduction and 4 main sections. In Sect. 2 we mentioned the basic facts and tools and mirror descent basics. Sect. 3 devoted to the proof of the optimal convergence rate of the mirror descent algorithm with time-varying step size rules for solving the classical unconstrained non-smooth optimization problem. In Sect. 4 the same analysis, that in Sect. 3, performed for more general optimization problems, namely the composite optimization problem. In Sect. 5 we present some numerical experiments which demonstrate the effectiveness of the proposed analysis.
@@ 2. Fundamentals
Let $Q\subset\mathbf{E}^{n}$ be a closed convex set, and $\psi:Q\longrightarrow\mathbb{R}$ be a proper closed differentiable and $\sigma$-strongly convex (called prox-function or distance generating function). The corresponding Bregman divergence is defined as
@@
(three points identity)[6] Suppose that $\psi:\mathbf{E}\longrightarrow(-\infty,\infty]$ is proper closed and convex, and differentiable over $\operatorname{dom}(\partial(\psi))$. Let $a,b\in\operatorname{dom}(\partial(\psi))$ and $c\in\operatorname{dom}(\psi)$. Then it holds
@@
3. Mirror descent algorithm
In this section, let us consider the following optimization problem
@@ 3. Mirror descent algorithm
Here and what follows, we denote the subdifferential of $f$ at $x$ by $\partial f(x)$, and the subgradient of $f$ at any point $x$ by $\nabla f(x)\in\partial f(x)$.
@@
0:  step sizes $\{\gamma_{k}\}_{k\geq 1}$, initial point $x^{1}\in Q$, number of iterations $N$.
@@
$\left(\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\right)\left[f\left(\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}\right)-f(x^{*})\right]\leq\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})-f(x^{*})\right).$
@@
$=\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+\sum_{k=2}^{N}\frac{1}{\gamma_{k}^{m+1}}V_{\psi}(x^{*},x^{k})-\sum_{k=2}^{N}\frac{1}{\gamma_{k-1}^{m+1}}V_{\psi}(x^{*},x^{k})$
$\;\;\;\;+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
@@
As a special case of Theorem 3.1, we can deduce the well-known sub-optimal convergence rate of mirror descent algorithm $O\left(\frac{\log(N)}{\sqrt{N}}\right)$, see [1, 2, 8, 7], with the following (adaptive and non-adaptive) time-varying step size rules
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with the time-varying step sizes given in (3.6), it satisfies the following sub-optimal convergence rate
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
@@
$f(\tilde{x})-f(x^{*})\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}.$
But
(3.9) $\sum_{k=1}^{N}\frac{1}{k}\leq 1+\log(N),\quad\text{and}\quad\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\geq 2\sqrt{N+1}-2,\quad\forall N\geq 1.$
@@
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}}.$
Also, from Theorem 3.1 we can obtain the optimal convergence rate of mirror descent algorithm $O\left(\frac{1}{\sqrt{N}}\right)$, with the time-varying step sizes given in (3.6).
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with the time-varying step sizes given in (3.6), it satisfies the following optimal convergence rate
@@
(3.12) $\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\leq 2\sqrt{N},\quad\forall N\geq 1.$
@@
$\leq\frac{1}{N}\frac{M_{f}}{\sqrt{2\sigma}}\left(\sqrt{N}V_{\psi}(x^{*},x^{1})+2\sqrt{N}\right)=\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
@@
By setting any $m>-1$ in (3.3) (in Theorem 3.1), the convergence rate of Algorithm 1 with the time-varying step sizes given in (3.6), will be optimal $O\left(\frac{1}{\sqrt{N}}\right)$ without presence of a $\log(N)$ factor. In comparison with the sub-optimal convergence rate (3.7), when $m>0$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points.
4. Mirror descent for composite problem
In this section, we will consider a more general problem than problem (3.1)
@@ 4. Mirror descent for composite problem
where $Q=\operatorname{dom}(h)\cap\operatorname{dom}(\partial\psi)$ and $f,h$ satisfy the following assumption
@@
    $f$ and $h$ are proper closed and convex.
@@
    $\operatorname{dom}(h)\subseteq\operatorname{int}(\operatorname{dom}(f))$.
@@
0:  step sizes $\{\gamma_{k}\}_{k\geq 1}$, initial point $x^{1}\in Q$, number of iterations $N$.
@@
    $\psi+\mathbb{I}_{\operatorname{dom}(\varphi)}$ be $\sigma$-strongly convex, where $\mathbb{I}_{A}$ is the indicator function of the set $A$.
@@
Then $a\in\operatorname{dom}(\partial(\psi))$, we have
@@
where $\hat{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ and $x^{*}$ is an optimal solution of (4.1).
@@
Now, from Lemma three points identity (2.2), with $a=x^{k+1},b=x^{k}$ and $c=u$, we get
@@
Now, in a similar way as in the proof of Theorem 3.1, we conclude that
@@
As a special case of Theorem 4.3, we can deduce the well-known sub-optimal convergence rate of mirror-C descent algorithm 2, namely $O\left(\frac{\log(N)}{N}\right)$, with the time-varying step sizes given in (3.6) (see [1, 2, 8, 7]).
@@
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with the time-varying step sizes given in (3.6), it satisfies the following sub-optimal convergence rate
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$.
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.13) we find
@@
Also, from Theorem 4.3 we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 2, with the time-varying step sizes given in (3.6).
@@
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with the time-varying step sizes given in (3.6), it satisfies the following optimal convergence rate
(4.14) $f(\overline{x})-f(x^{*})\leq\left(\frac{M_{f}V_{\psi}(x^{*},x^{1})+2M_{f}}{\sqrt{2\sigma}}\right)\frac{1}{\sqrt{N}}=O\left(\frac{1}{\sqrt{N}}\right),$
@@
$F(\overline{x})-F(x^{*})$ $\leq\frac{1}{N}\left(\frac{M_{f}\sqrt{N}V_{\psi}(x^{*},x^{1})}{\sqrt{2\sigma}}+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\right)$

### 25 Jan 2024 v3

# Optimal Convergence Rate for Mirror Descent Methods with special Time-Varying Step Sizes Rules

@@ Abstract.
In this paper, the optimal convergence rate $O\left(N^{-1/2}\right)$ (where $N$ is the total number of iterations performed by the algorithm), without the presence of a logarithmic factor, is proved for mirror descent algorithms with special time-varying step sizes, for solving classical constrained non-smooth problems, problems with the composite model and problems with non-smooth functional (inequality types) constraints. The proven result is an improvement on the well-known rate $O\left(\log(N)N^{-1/2}\right)$ for the mirror descent algorithms with the time-varying step sizes under consideration. It was studied a new weighting scheme assigns smaller weights to the initial points and larger weights to the most recent points. This scheme improves the convergence rate of the considered mirror descent methods, which in the conducted numerical experiments outperform the other methods providing a better solution in all the considered test problems.
@@
Convex optimization, Non-smooth problem, Composite problem, Problem with functional constraints, Mirror descent, Mirror-C descent, Optimal convergence rate, Sub-optimal convergence rate, Time-varying step size.
@@ 1. Introduction
The first method for unconstrained minimization of non-smooth convex function was proposed with fixed step size in [31]. In the next years, there were developed several strategies for choosing the step sizes [29]. Among them, are fixed step length, non-summable diminishing step, square summable but not summable step [11], Polyak step [29], and many others (see Table 1, below). The optimal convergence rate of the subgradient method is known as $O\left(N^{-1/2}\right)$, where $N$ is the total number of iterations performed by the algorithm [4, 12, 20, 25]. This rate was proved by using many rules of the step size. One of them is fixed (constant), namely $\gamma_{k}=O(1)/\sqrt{N}$ for $k=1,\ldots,N$ (with special chosen of the constant $O(1)$), which minimizes the upper bound of the difference between the best value of the objective function attained (i.e., the best value observed during the minimization process) and the optimal value. But most of the time it is preferable to use the time-varying step size since the constant step size rule requires the a priori knowledge of the total number of iterations employed by the method. From the mentioned fixed steps, the inspiration was to take $\gamma_{k}=O(1)/\sqrt{k}$ for $k=1,\ldots,N$, but the convergence rate, with this rule, became sub-optimal $O\left(\log(N)N^{-1/2}\right)$. Recently in [39], for the projected subgradient method it was proved the optimal convergence rate using the previously mentioned time-varying step size with a new weighting scheme for the generated points each iteration of the algorithm.
@@
The next important advancement in this area is related to the development of the mirror descent method which originated in [23, 24] and was later analyzed in [5]. It is considered as the non-Euclidean extension of standard subgradient methods. This method is used in many applications, see [8, 18, 26, 27, 35] and references therein. The standard subgradient methods employ the Euclidean distance function with a suitable step size in the projection step. Mirror descent extends the standard projected subgradient methods by employing a nonlinear distance function with an optimal step size in the nonlinear projection step [22]. The mirror descent method is also applicable to optimization problems in Banach spaces where gradient descent is not [14]. An extension of the mirror descent method for constrained problems was proposed in [6, 24]. Usually, the step size and stopping criterion for the mirror descent method require knowing the Lipschitz constant of the objective function and constraints, if any. Adaptive step sizes, which do not require this information, are considered for unconstrained problems in [7], and for constrained problems in [6].
In addition to the previously mentioned problems, there is a more general class of problems, which is non-smooth optimization problems with non-smooth functional (inequality types) constraints (see problem (5.1) or its equivalence (5.2)). This class of problems arises and attracts widespread interest in many areas of modern large-scale optimization and its applications [9, 30]. On continuous optimization with functional constraints, there is a long history of studies. The monographs in this area include [10, 28]. Some of the works on first-order methods for convex optimization problems with convex functional constraints include [3, 17, 21, 32, 33, 34, 37] for the deterministic setting and [1, 2, 19, 38] for the stochastic setting.
@@
As we pointed out the mirror descent algorithm was proposed and analyzed a long time ago for different step size rules. However, the optimal convergence rate $O(N^{-1/2})$ of the algorithm was first proved, as for the subgradient method, with the fixed step size $\gamma_{k}=O(1)/\sqrt{N}$ for $k=1,\ldots,N$. But, when the step sizes are time-varying, namely $\gamma_{k}=O(1)/\sqrt{k}$, it was proved the sub-optimal convergence rate $O(\log(N)N^{-1/2})$, see [4, 12, 20, 25] for details. Although the constant step size rule is relatively easy to analyze but has the disadvantage of requiring the a priori knowledge of the total number of iterations employed by the method. In practical situations, the number of iterations is not fixed a priori, and stopping criteria different than merely fixing the total number of iterations is usually imposed. This is why dynamic, or time-varying (i.e., non-constant) step size rules are important [12].
@@
In this paper, we extend the results of [39] to the mirror descent methods. We proved the optimal convergence rate, without the presence of a logarithmic factor, for mirror descent methods with special time-varying step sizes, for solving classical constrained non-smooth problems (see Sect. 3), constrained composite problems (see Sect. 4), and problems with functional constraints (see Sect. 5). We demonstrate the efficiency of the theoretical studied results by some conducted experiments for some constrained (with and without functional constraints) optimization problems, such as best approximation problem, Fermat-Torricelli-Steiner problem, smallest covering ball problem and the maximum of a finite collection of linear functions.
@@
The paper consists of an introduction and 5 main sections. In Sect. 2 we mentioned the basic facts and tools and mirror descent basics. Sect. 3 devoted to the proof of the optimal convergence rate of the mirror descent algorithm with time-varying step size rules for solving the classical constrained non-smooth optimization problem. In Sect. 4 the same analysis, that in Sect. 3, performed for more general optimization problems, namely the constrained composite optimization problem. Section 5 is devoted to the problem with functional constraints, which is a more general class of non-smooth optimization problems. In Sect. 6 we present some numerical experiments which demonstrate the effectiveness of the proposed algorithms and analysis.
@@ 2. Fundamentals
Let $Q\subset\mathbf{E}^{n}$ be a compact convex set, and $\psi:Q\longrightarrow\mathbb{R}$ be a proper closed differentiable and $\sigma$-strongly convex (called prox-function or distance generating function) with $\sigma>0$. The corresponding Bregman divergence is defined as
@@
(three points identity) [13] Suppose that $\psi:\mathbf{E}\longrightarrow(-\infty,\infty]$ is proper closed, convex, and differentiable over $\operatorname{dom}(\partial(\psi))$. Let $a,b\in\operatorname{dom}(\partial(\psi))$ and $c\in\operatorname{dom}(\psi)$. Then it holds
@@
In what follows, we denote the subdifferential of $f$ at $x$ by $\partial f(x)$, and the subgradient of $f$ at any point $x$ by $\nabla f(x)\in\partial f(x)$.
3. Mirror descent method for constrained problem
In this section, we consider the following constrained optimization problem
@@ 3. Mirror descent method for constrained problem
For problem (3.1), we use Algorithm 1, under consideration
(3.3) $\max_{x\in Q}V_{\psi}(x^{*},x)\leq V_{\psi}(x^{*},x^{1})<\infty,$
where $x^{1}\in Q$ is a chosen (dependently on $Q$) initial point and $x^{*}$ is an optimal solution of (3.1) nearest to $x^{1}$.
@@
0:  step sizes $\{\gamma_{k}\}_{k\geq 1}$, initial point $x^{1}\in Q$ s.t. (3.3) holds, number of iterations $N$.
@@
$\qquad\qquad\qquad\qquad\qquad\leq\sum_{k=1}^{N}\frac{1}{\gamma_{k}^{m}}\left(f(x^{k})-f(x^{*})\right).$
@@
Due to (3.3), we find that $V_{\psi}(x*,x^{k})\leq V_{\psi}(x*,x^{1}),\,\forall k=2,\ldots,N$. Hence we get
$\left(\sum_{k=1}^{N}\gamma_{k}^{-m}\right)\left(f(\hat{x})-f(x^{*})\right)\leq\frac{1}{\gamma_{1}^{m+1}}V_{\psi}(x^{*},x^{1})+\sum_{k=2}^{N}\frac{1}{\gamma_{k}^{m+1}}V_{\psi}(x^{*},x^{k})$
$\;\;\;\;-\sum_{k=2}^{N}\frac{1}{\gamma_{k-1}^{m+1}}V_{\psi}(x^{*},x^{k})+\frac{1}{2\sigma}\sum_{k=1}^{N}\frac{\|\nabla f(x^{k})\|_{*}^{2}}{\gamma_{k}^{m-1}}$
@@
As a special case of Theorem 3.1, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate of mirror descent algorithm $O\left(\frac{\log(N)}{\sqrt{N}}\right)$, see [4, 12, 20, 25], with the following (adaptive and non-adaptive) time-varying step size rules
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with $m=-1$, and the time-varying step sizes given in (3.7), it satisfies the following sub-optimal convergence rate
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}}\sum_{k=1}^{N}\gamma_{k}x^{k}$.
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}}\sum_{k=1}^{N}\gamma_{k}x^{k}$.
@@
$f(\tilde{x})-f(x^{*})\leq\frac{M_{f}}{\sqrt{2\sigma}}\frac{V_{\psi}(x^{*},x^{1})+\sum_{k=1}^{N}\frac{1}{k}}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$. But
$\sum_{k=1}^{N}\frac{1}{k}\leq 1+\log(N),\quad\text{and}\quad\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\geq 2\sqrt{N+1}-2,\quad\forall N\geq 1.$
@@
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}}\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}x^{k}$. ∎
Also, from Theorem 3.1, with a special value of the parameter $m$, we can obtain the optimal convergence rate of mirror descent algorithm $O\left(\frac{1}{\sqrt{N}}\right)$, with the time-varying step sizes given in (3.7).
@@
Let $f$ be an $M_{f}$-Lipschitz convex function, and assume $\|\nabla f(x)\|_{*}\leq M_{f},\forall x\in Q$. Then for problem (3.1), by Algorithm 1, with $m=0$, and the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
@@
$\sum_{k=1}^{N}\frac{1}{\sqrt{k}}\leq 2\sqrt{N},\quad\forall N\geq 1.$
@@
$=\frac{M_{f}(2+V_{\psi}(x^{*},x^{1}))}{\sqrt{2\sigma}}\frac{1}{\sqrt{N}}.$
@@
By setting any $m>-1$ in (3.4) (in Theorem 3.1), the convergence rate of Algorithm 1 with the time-varying step sizes given in (3.7), will be optimal $O\left(\frac{1}{\sqrt{N}}\right)$ without presence of a $\log(N)$ factor. In comparison with the sub-optimal convergence rate (3.8), when $m>0$, the weighting scheme $\frac{1}{\sum_{k=1}^{N}\gamma_{k}^{-m}}\sum_{k=1}^{N}\gamma_{k}^{-m}x^{k}$ assigns smaller weights to the initial points and larger weights to the most recent points that generated by Algorithm 1. This fact will be shown in numerical experiments.
4. Mirror descent method for constrained composite problem
In this section, we will consider a more general problem than problem (3.1), namely
@@ 4. Mirror descent method for constrained composite problem
where $Q=\operatorname{dom}(h)\cap\operatorname{dom}(\partial\psi),f$ and $h$ satisfy the following assumption
@@
    $f$ and $h$ are proper closed and convex;
@@
    $\operatorname{dom}(h)\subseteq\operatorname{int}(\operatorname{dom}(f))$;
@@
0:  step sizes $\{\gamma_{k}\}_{k\geq 1}$, initial point $x^{1}\in Q$ s.t. (3.3) holds, number of iterations $N$.
@@
    $\psi+\mathbb{I}_{\operatorname{dom}(\varphi)}$ be a $\sigma$-strongly convex, where $\mathbb{I}_{A}$ is the indicator function of the set $A$.
@@
Then we have
@@
Now, from three points identity (2.2), with $a=x^{k+1},b=x^{k}$ and $c=u$, we get
@@
Now, in a similar way as in the proof of Theorem 3.1, we conclude
@@
As a special case of Theorem 4.3, with a special value of the parameter $m$, we can deduce the well-known sub-optimal convergence rate of the mirror-C descent method (Algorithm 2), namely $O\left(\frac{\log(N)}{N}\right)$, with the time-varying step sizes given in (3.7) (see [4, 12, 20, 25]).
@@
Suppose that Assumption 4.1 holds and $h$ is a non-negative function. Then for problem (4.1), by Algorithm 2, with $m=-1$, and the time-varying step sizes given in (3.7), it satisfies the following sub-optimal convergence rate
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}}\sum_{k=1}^{N}\gamma_{k}x^{k}$.
@@
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\gamma_{k}}\sum_{k=1}^{N}\gamma_{k}x^{k}$.
@@
When $\gamma_{k}=\frac{\sqrt{2\sigma}}{M_{f}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.13), with $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\frac{1}{\sqrt{k}}}\sum_{k=1}^{N}\frac{1}{\sqrt{k}}x^{k}$, we find
@@
Case 2 (adaptive rule). When $\gamma_{k}=\frac{\sqrt{2\sigma}}{\|\nabla f(x^{k})\|_{*}\sqrt{k}},\,k=1,2,\ldots,N$, and since $\|\nabla f(x^{k})\|_{*}\leq M_{f}$, then by substitution in (4.13), we find
@@
$\leq\frac{M_{f}}{\sqrt{\sigma}}\frac{\frac{\sqrt{2\sigma}}{M_{f}}h(x^{1})+V_{\psi}(x^{*},x^{1})+1+\log(N)}{\sqrt{N}},$
where $\tilde{x}=\frac{1}{\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}}\sum_{k=1}^{N}\left(\|\nabla f(x^{k})\|_{*}\sqrt{k}\right)^{-1}x^{k}$. ∎
Also, from Theorem 4.3, with a special value of the parameter $m$, we can obtain the optimal convergence rate $O\left(\frac{1}{\sqrt{N}}\right)$ of Algorithm 2, with the time-varying step sizes given in (3.7).
@@
Suppose that Assumption 4.1 hold and $h$ is a non-negative function, with $h(x^{1})=0$. Then for problem (4.1), by Algorithm 2, with $m=0$, and the time-varying step sizes given in (3.7), it satisfies the following optimal convergence rate
$F(\overline{x})-F(x^{*})\leq\left(\frac{M_{f}V_{\psi}(x^{*},x^{1})+2M_{f}}{\sqrt{2\sigma}}\right)\frac{1}{\sqrt{N}}=O\left(\frac{1}{\sqrt{N}}\right),$